The Sullivan Co. needs to raise $65.8 million to finance its expansion into new markets. The company will sell new shares of equity via a general cash offering to raise the needed funds. The offer price is $58 per share and the company's underwriters charge a spread of 8 percent. How many shares need to be sold?
Answer #1

Required proceeds net of spread=65,800,000/(1-spread)

=65,800,000/(1-0.08)

=$71521739.1

Hence shares to be sold=71521739.1/58

=1233133 shares(Approx).